A monopolist with the cost function C=1/2Q^2 faces a demand curve Q=12-P,



A. What will be his equilibrium price and quantity?



B. If for some reason the firm behaves as if it were in a perfectly competitive industry, what will equilibrium price and quantity?



C. How much money will the firm require to forgo monopoly profits and behave competitively instead?




1
Expert's answer
2022-02-21T13:20:52-0500

A) Given that "C=\frac{1}{2}Q^2" and "Q=12-P"

To find the price, P, we take the inverse of Q, so that

"P=12-Q"


The total revenue,"TR=P\u00d7Q"

"TR=(12-Q)Q=12Q-Q^2"

And

"MR=TR'=12-2Q"


"MC=TC'=Q"

At equilibrium, "MR=MC"

"\therefore 12-2Q=Q"

"Q^*=4"

"P^*=12-4=8"


B) In a perfectly competitive industry, "P=MC" , which implies that the firm would charge a lower price but now a sells a higher quantity.

"\therefore8=Q^*"

And "P^*=12-8=4"


C)

The monopolist's profit would have been

"\pi=TR-TC"

At Q* = 4

"\pi=12(4)-(4^2)-\frac{1}{2}(4^2)"

"\pi=24"


The monopolist's profit, if it behaves like a perfectly competitive industry, would be

"\pi=12(8)-(8^2)-\frac{1}{2}(8^2)"

"\pi=0"



Therefore, the monopolist would be giving up all of it's profit (24) to behave like a perfectly competitive industry.